# Exact output tracking for the one-dimensional heat equation and applications to the interpolation problem in Gevrey classes of order 2

### v1

@@ 4 Interpolation in Gevrey classes
We begin by proving the result in case $R=1/\sqrt{2}$ and $\gamma=0$. This case is special because the system ‘‘Dirichlet-to-Dirichlet’’ (39) has its trackable subspace satisfying
@@
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Dir}}=\{f\in\operatorname{Hol}(\Omega):f\mbox{ even},\quad f\in L^{2}(\Omega)\}.$
@@
By cutoff arguments, there is no loss of generality in assuming that $I$ is bounded (see [reachable_martin, Lemma 3.7]). By translation we may further assume that $\inf I=0$. Then, on the one hand, take $(a_{k})$ in $\mathcal{B}_{t_{0}}\mathcal{G}_{2,1/\sqrt{2},0}(I)$ and fix $\varphi$ a pre-image. By Corollary 3 the function $\varphi$ is in $\mathcal{Y}_{\operatorname{Dir}-\operatorname{Dir}}(0,t_{0})$ and thus satisfies
@@
By definition of $\mathcal{R}_{\operatorname{Dir}-\operatorname{Dir}}$ and the properties of the reachable space of the heat equation, there is $u\in L^{2}(0,t_{0})$ such that the solution $z$ of (39) with control $u$ reaches the state defined by the power series in (45), in time $t_{0}$. We extend the control $u$ by 0 on $(t_{0},\infty)$, which extends $z$ for any positive time, what shows that $(a_{k})\in\mathcal{B}_{t_{0}}\mathcal{G}_{2,1/\sqrt{2},0}(I)$. This completes the proof of (43).
@@ Appendix A Appendix: more on the interpolation problem
In this Appendix, we provide a detailed analysis of the interpolation problem in Gevrey classes $C^{\infty}(k^{ks})$, mostly when $s>1$, notably by completing the proof of [mitjagin, Theorems 1 and 1a], which is only sketched in [mitjagin] (see also [mitjagin_english]).
Let us first clarify Remark 5: in [reachable_martin, Theorem 3.2], it is claimed that for every sequence $(a_{k})$ and positive numbers $A,\epsilon$, the following holds:
$|a_{k}|\lesssim A^{k}(2k)!\Longrightarrow(a_{k})\in\mathcal{B}C^{\infty}((2k)!,\rho_{MRR}A+\epsilon),$ (46)
with the loss factor
$\rho_{MRR}:=\exp\left(\frac{1}{e}\right)\approx 1.445<2=\Gamma_{2}.$
This contradicts Proposition 4, which requires the loss factor in (46) to be at least $2$. This also contradicts [mitjagin, Theorem 1a] (see also (17)), which states that the loss factor $\rho$ must satisfy $\rho\geq\Gamma_{2}=2$. We believe that, unfortunately, the proof of [reachable_martin, Theorem 3.2] is flawed. Indeed, to go further into the details of the proof in [reachable_martin], the essence of their Theorem 3.2 lies in their Proposition 3.6. To prove this last Proposition, the authors build $f(x)=\sum_{p}d_{p}\varphi_{p}(x)x^{p}/p!$ where each $\varphi_{p}$ is built according to their Corollary 3.5. The constant $C$ appearing in their Corollary 3.5 implicitly depends on $\delta$ and on the sequence $(a_{k})$. However, in the proof of their Proposition 3.6, this dependency is not taken into account, and the constant $C$ appearing in the estimate of $\varphi_{p}^{(j)}(x)$ page 195 a priori depends at least on $p$, which prevents one from concluding as the authors claim. We believe this is a minor inaccuracy, since the proof of [reachable_martin, Theorem 3.2] can easily be corrected to yield a valid result, with an increased (and probably suboptimal) constant $\rho$, see [petzsche, Theorem 3.6]. Moreover, this does not detract from the importance of the article [reachable_martin], which is the first to successfully describe the reachable states of the heat equation using holomorphic functions.
@@ A.1 Generalities
$\infty<a<b<\infty,\quad c\in[a,b],\quad A>0,\quad 1\leq\rho<\infty,\quad 1<s<\infty.$
We say that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss $\rho$ if, for every sequence of complex numbers $(a_{n})$ such that
@@
$\exists C^{\prime}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[a,b],\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns}.$
The loss factor $\rho$ is the quantity we aim to estimate. The above problem a priori depends on the parameters $(a,b,c,A,s)$. We claim that actually it does not depend on $a,b,c,A$ when $s$ is fixed, in the sense that changing $a,b,c$ and $A$ does not affect the loss factor $\rho$.
@@
Fix $s>1$ and assume that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss $\rho$. Then, the interpolation problem associated with $(a^{\prime},b^{\prime},c^{\prime},A^{\prime},s)$ is solvable with the same loss factor $\rho$.
@@
Assume that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss factor $\rho$. We show that for every $(a^{\prime},b^{\prime},c^{\prime},A^{\prime},s)$ as above, the associated problem is solvable with loss $\rho$. Clearly, for any translation $\tau\in\mathbb{R}$, the problem associated with $(a+\tau,b+\tau,c+\tau,A,s)$ is solvable with the same loss factor. We may therefore assume that $a=a^{\prime}=0$. Let us then use a dilatation: if $\varphi\in C^{\infty}[0,b]$ is such that
$|\varphi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns},$
@@
$\varphi_{\lambda}\in C^{\infty}[0,b/\lambda],\quad\varphi_{\lambda}^{(n)}(t)=\lambda^{n}\varphi^{(n)}(t),\quad|\varphi_{\lambda}^{(n)}(t)|\leq C^{\prime}(\rho A\lambda+\epsilon\lambda)^{n}n^{ns}.$
Thus the problem of parameters $(0,b/\lambda,c/\lambda,A\lambda,s)$ is solved with loss $\rho$. We can therefore assume that $b=b^{\prime}=1$.
We then claim that if the problem of parameters $(0,1,c,A,s)$ is solvable with loss $\rho$ for $0<c<1$, then for every $c^{\prime}\in(0,1)$ the problem of parameters $(0,1,c^{\prime},A,s)$ is solvable with loss $\rho$. For, let $\epsilon>0$ and $(a_{n})$ be any sequence such that
@@
$\varphi^{(n)}(c)=a_{n},\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns}.$
@@
$|\psi^{(n)}(t)|\leq C^{\prime\prime}(\rho A+\epsilon)^{n}n^{ns},$
(see [reachable_martin, Lemma 3.7]). Moreover, since $\chi$ is 0 outside of $(c^{\prime}-c,1+c^{\prime}-c)$, $\psi$ is well-defined on $\mathbb{R}$ (so it is also defined on $[0,1]$), and has the same derivatives at $c^{\prime}$ than $\varphi$ has at $c$. Thus the problem of parameters $(0,1,c^{\prime},A,s)$ is solvable, with the same loss $\rho$.
Next we show that the cases $c=0$ or $c=1$ are equivalent. Assume for instance that the problem of parameters $(0,1,0,A,s)$ is solvable with loss $\rho$. For any sequence $(a_{n})$ such that
@@
$\varphi^{(n)}(c)=(-1)^{n}a_{n},\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns}.$
@@
Assume now that the problem of parameters $(0,1,0,A,s)$ is solvable with loss $\rho$. We show that, for every $0<c<1$ the problem of parameters $(0,1,c,A,s)$ is solvable with loss $\rho$. In fact we will show that (1,1,0,A,s)(-1,1,0,A,s) is solvable with loss $\rho$, which is easily seen as a stronger statement, by translation and dilatation arguments. We let $(a_{n})$ be a sequence with
@@
We introduce the ‘‘odd’’ and ‘‘even’’ sequences
@@
$|\varphi^{(n)}(t)|,|\psi^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns},\quad\varphi^{(n)}(0)=a_{n}^{o},\quad\psi^{(n)}(0)=a_{n}^{e}.$
@@
$|\tilde{\psi}^{(n)}(t)|,|\tilde{\varphi}^{(n)}(t)|\leq C^{\prime}(\rho A+\epsilon)^{n}n^{ns}.$
@@
We can thus put $c=c^{\prime}=0$, hence we are left to show that if the problem $(0,1,0,A,s)$ is solvable with loss $\rho$, then, for every $A^{\prime}>0$, the problem $(0,1,0,A^{\prime},s)$ is solvable with loss $\rho$. Assume that the problem associated with $(0,1,0,A,s)$ is solvable with loss $\rho$, using the dilatation $\varphi_{\lambda}$ we see that for every $\lambda>0$, the problem associated with $(0,1/\lambda,0,A\lambda,s)$ is solvable with loss $\rho$. We distinguish two cases: if $\lambda\leq 1$, then the problem associated with $(0,1,0,\rho,A\lambda,s)$ is trivially solvable. If on the contrary $\lambda>1$, using a cutoff that is Gevrey of order $1<r<s$, constant to $1$ near 0 and flat at $1/\lambda$, we see that the problem associated with $(0,1,0,\rho,A\lambda,s)$ is also solvable with loss $\rho$, as in the case of the equivalence for the parameter $c$. Since $\lambda>0$ is arbitrary, this finishes the proof. ∎
Therefore, the only relevant numbers are $\rho$ and $s$.
@@ A.2 A proof of Mitjagin’s Theorem
    We have the estimation on the Taylor coefficients of $\Phi$
@@
    where, as recalled, $r=|s|$. In view of the definition of the $\|\cdot\|_{C^{\infty}_{\beta B}}$-norm, the estimate (47) thus boils down to proving that
@@
    i.e. it is of order $1$ and of type 0 on $\mathbb{C}$. Note however that on the real axis we have the better estimate
@@
    Let $0<\rho<1$ and $M,K\geq 0$. Assume that $g\in\operatorname{Hol}(\mathbb{C}_{+})\cap C^{0}(\mathbb{C}_{+}\cup i\mathbb{R})$ is of order 1 with minimal type 0, together with
@@
    Let us show this version of the Phragmén-Lindelöf principle. We will use the following ‘‘standard’’ version of the Phragmén-Lindelöf principle, see [K1, §III.C]
    Let $\alpha\geq 1/2$, $\Omega=\{re^{i\theta}:r>0,\quad|\theta|<\frac{\pi}{2\alpha}\}$, and $g\in\operatorname{Hol}(\Omega)\cap C^{0}(\overline{\Omega})$ of order $\rho<\alpha$, such that
@@
    $g_{\kappa}(z)=\exp\left(-\frac{K}{\cos(\pi\rho/2)}z^{\rho}-\kappa z\right)g(z)\in\operatorname{Hol}(\mathbb{C}_{+})\cap C^{0}(\mathbb{C}_{+}\cup i\mathbb{R}).$
    We apply the ‘‘standard’’ Phragmén-Lindelöf’s method quoted above to the function $g_{\kappa}$ on the domain $\Omega:=\{x+iy:x,y>0\}$. Note that the quoted result is invariant by rotation, hence we may see $\Omega$ as a sector with angle $\pi/2$, hence the parameter $\alpha$ is set to $2$. On $\Omega$ the function $g_{\kappa}$ is of order $\rho=1$, which is strictly less than $\alpha$. Moreover, on the boundary of $\Omega$ lying on the imaginary axis there holds
@@
    We take $\eta=\kappa$ so that for every $x>0$
@@
    $|f_{n}|=\left|\frac{\Phi^{(n)}(0)}{n!}\right|\leq\frac{1}{r^{n}}\max_{|s|=r}|\Phi(s)|\leq\frac{1}{r^{n}}\|L\|C_{1}\exp\left[\frac{\beta}{e\cos(\pi/2\beta)}\left(\frac{1+\delta}{B}r\right)^{1/\beta}\right],$
@@
    $|f_{n}|\leq\|L\|C_{1}\left(\frac{e\mu}{n\beta}\right)^{n\beta}=C_{1}\|f\|\left(\frac{1+\delta}{B\cos^{\beta}\frac{\pi}{2\beta}}\right)^{n}n^{-n\beta}.$
@@
where the inclusions are continuous and dense [^6] That the first inclusion is bounded is trivial. That the second inclusion is bounded can be shown using the Sobolev embeding $H^{1}(-1,1)\subset C[-1,1]$. That the inclusions are dense is shown by mollifying with a kernel in $C^{\infty}_{\gamma C}$ for some $1<\gamma<\beta$ and $C>0$.. We will need a similar construction for our spaces of sequences. For a sequence $(w_{n})$ of positive numbers we define $K(w_{n})$ as the set of these sequences $a=(a_{n})$ such that
@@
which is what we aim to show until the end of the proof. From the density of the set of finitely supported sequences in $Y$, which we denote by $\mathbb{C}^{(\mathbb{N})}$, it is enough to show (54) when $a\in\mathbb{C}^{(\mathbb{N})}$. For such $a$ consider the concentrated functional
$f:=\sum_{n=0}^{\infty}a_{n}\delta^{(n)}.$
@@
$\exists c>0,\quad\forall a\in\mathbb{C}^{(\mathbb{N})},\quad\sum_{n=0}^{\infty}|B^{n}n^{n\beta}a_{n}|^{2}\leq c\|f\|_{X^{\prime}}^{2}.$
We then invoke Lemma 28 for the functional $f$. To this end, observe that
@@
so that Lemma 28 applies to $f$. Then put
$\Phi(s)=\sum_{n=0}^{\infty}f_{n}s^{n},\quad f_{n}=a_{n},$
@@
$|a_{n}|\leq C_{1}\|f\|_{X^{\prime}}\left(\frac{1+\delta}{\Gamma_{\beta}B+\epsilon/2}\Gamma_{\beta}\right)^{n}n^{-n\beta}.$
We deduce
$|B^{n}n^{n\beta}a_{n}|\leq C_{1}\|f\|_{X^{\prime}}\left(\frac{1+\delta}{1+\frac{\epsilon}{2\Gamma_{\beta}B}}\right)^{n},$
@@
Douglas’s lemma yields a slightly better conclusion than the inclusion (54): there exists a bounded operator $\mathfrak{R}:\ell^{2}_{\beta B}\rightarrow L^{2}_{\beta,\Gamma_{\beta}B+\epsilon}$ such that $i=\mathcal{B}\mathfrak{R}$. In other words, the interpolation problem in the Gevrey class $C^{\infty}(k^{ks})$ can be solved continuously and linearly. Compare with [mitjagin_bases, Proposition 21, §7, p. 119].

### v2

@@ 4 Interpolation in Gevrey classes of order 22
We begin by proving the result in case $R=1/\sqrt{2}$ and $\gamma=0$. This case is special because the system “Dirichlet-to-Dirichlet” (41) has its trackable subspace satisfying
@@
$\mathcal{R}_{\operatorname{Dir}-\operatorname{Dir}}=\{f\in\mathcal{H}(\Omega):f\mbox{ even},\quad f\in L^{2}(\Omega)\}.$ (45)
@@
By cutoff arguments, there is no loss of generality in assuming that $I$ is bounded (see [reachable_martin, Lemma 3.7]). By translation we may further assume that $\inf I=0$. Then, on the one hand, take $(a_{k})$ in $\mathcal{B}_{t_{0}}\mathcal{G}_{2,1/\sqrt{2},0}(I)$ and fix $\varphi$ a pre-image. Owing to Corollary 3 the function $\varphi$ is in $\mathcal{Y}_{\operatorname{Dir}-\operatorname{Dir}}(0,t_{0})$ and thus satisfies
@@
By definition of the reachable space and (45), there is $u\in L^{2}(0,t_{0})$ such that the solution $z$ of (41) with control $u$ reaches the state defined by the power series in (48), in time $t_{0}$. We extend the control $u$ by 0 on $(t_{0},\infty)$, which extends $z$ for any positive times, what shows that $(a_{k})\in\mathcal{B}_{t_{0}}\mathcal{G}_{2,1/\sqrt{2},0}(I)$. This completes the proof of (46).
@@
Proof of Corollary 10.
We fix $(a_{n})$ an arbitrary sequence of complex numbers and we introduce the quantity
$\mathfrak{R}_{a}:=\sup\{R>0:\exists\varphi\in\mathcal{G}_{2,R,0}(-1,1),\quad\forall n\in\mathbb{N},\quad\varphi^{(n)}(0)=a_{n}\}.$
From Proposition 9 we clearly have
$\mathfrak{R}_{a}=\sup\left\{R\geq 0:\sum_{k=0}^{\infty}a_{k}\frac{(\sqrt{2}R\zeta)^{2k}}{(2k)!}\in A^{2}(\Omega)\right\},$
hence we are left to show that $\mathfrak{R}_{a}=R_{a}$. On the one hand, for all $0<R^{\prime}<R$ we have $G^{2,R}[-1,1]\subset\mathcal{G}_{2,R^{\prime},0}(-1,1)$, hence $R_{a}\leq\mathfrak{R}_{a}$. On the other hand, let $0<R<\mathfrak{R}_{a}$ and $\varphi\in\mathcal{G}_{2,R,0}(-1,1)$ interpolating $(a_{n})$. From [reachable_martin, Lemma 3.7] we can further assume that $\varphi$ is flat at $\pm 1$, so that in particular the Poincaré inequality applies to $\varphi$ and and all its successive derivatives. We easily deduce that for all $0<R^{\prime}<R$ there holds $\varphi\in G^{2,R^{\prime}}[-1,1]$, hence $R^{\prime}\leq R_{a}$. We deduce that $\mathfrak{R}_{a}\leq R_{a}$, which concludes the proof. ∎
In Proposition 4 we have given an example of a sequence $(a_{k})$ whose Gevrey radius is $1$, but which cannot be interpolated by a function that is Gevrey of radius $1/\sqrt{2}$. Corollary 10 allows to obtain the other extreme case: the sequence
$a_{k}:=\frac{(2k)!2^{k}i^{k}}{\sqrt{k}},\quad a_{0}:=0,$
can be interpolated by a function which is Gevrey of order $2$ and radius $1/\sqrt{2}-\epsilon$, for all $\epsilon>0$ (see (38)).
@@ Appendix A Appendix: more on the interpolation problem
In this Appendix, we provide a detailed analysis of the interpolation problem in Gevrey classes of order $s>1$, notably by completing the proof of Theorems 1 and 1a of [mitjagin], which is only sketched in the cited paper.
Let us first clarify Remark 5: in [reachable_martin, Theorem 3.2], it is claimed that for every sequence $(a_{n})$ such that
$|a_{n}|\lesssim\frac{(2n)!}{R_{0}^{2n}},$
and for every $\epsilon>0$, the sequence $(a_{n})$ can be interpolated by a function $\varphi\in C^{\infty}(\mathbb{R})$ such that
$|\varphi^{(n)}(t)|\lesssim\frac{(2n)!}{(\rho^{MRR}R_{0}-\epsilon)^{2n}},$
where
$\rho^{MRR}:=\exp\left(-\frac{1}{2e}\right)\approx 0.832>0.707\approx\frac{1}{\sqrt{2}}=\rho_{2}.$
This contradicts Proposition 4 and [mitjagin, Theorem 1a], which impose the loss factor $\rho^{MRR}$ to be $\geq 1/\sqrt{2}$.
We believe that, unfortunately, the proof of [reachable_martin, Theorem 3.2] is flawed. Indeed, to go further into the details of the proof in [reachable_martin], the essence of their Theorem 3.2 lies in their Proposition 3.6. To prove this last Proposition, the authors build $f(x)=\sum_{p}d_{p}\varphi_{p}(x)x^{p}/p!$ where each $\varphi_{p}$ is built according to their Corollary 3.5. The constant $C$ appearing in their Corollary 3.5 implicitly depends on $\delta$ and on the sequence $(a_{k})$. However, in the proof of their Proposition 3.6, this dependency is not taken into account, and the constant $C$ appearing in the estimate of $\varphi_{p}^{(j)}(x)$ page 195 a priori depends at least on $p$, which prevents one from concluding as the authors claim.
Note further that [reachable_martin, Proposition 3.6] may be used to show that the interpolation problem can be solved in Gevrey classes of order $s>1$, with loss factor
$\rho_{s}^{MRR}:=\exp\left(-\frac{1}{es}\right).$
As explained above, we believe that the proof of [reachable_martin, Proposition 3.6] is flawed. In fact, for $1<s<3$ we have $\rho_{s}^{MRR}>\rho_{s}$, hence a contradiction with Mitjagin’s results. For $s>4$ we have $\rho_{s}^{MRR}<\rho_{s}$, hence $\rho_{s}^{MRR}$ is a valid loss factor, although sub-optimal.
We believe this is a minor inaccuracy. Indeed, the result itself of [reachable_martin, Theorem 3.2] might be corrected to yield a valid result, by considering the loss factor of [mitjagin, Theorem 1], whereas the proof of [reachable_martin, Theorem 3.2] might be corrected to yield a valid result, probably with a suboptimal loss factor, see [petzsche, Theorem 3.6]. As far as we are concerned, whenever Theorem 3.2 or Proposition 3.6 from [reachable_martin] is used, the resulting assertion can be corrected by merely modifying a numerical constant. Moreover, this does not detract from the importance of the article [reachable_martin], which is the first to successfully describe the reachable states of the heat equation using holomorphic functions.
@@ A.1 Generalities
In this subsection we study the interpolation problem in a rather general setting. In order to remain close to Mitjagin’s notations we change our convention for Gevrey functions. This is a minor modification and we shall explain how to relate the loss factors between the two conventions.
@@
$\infty<a<b<\infty,\quad c\in[a,b],\quad A>0,\quad 1\leq\Gamma<\infty,\quad 1<s<\infty.$
We say that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss $\Gamma$ if, for every sequence of complex numbers $(a_{n})$ such that
@@
$\exists C^{\prime}>0,\quad\forall n\in\mathbb{N},\quad\forall t\in[a,b],\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns}.$
If $\Gamma_{s}$ stands for the optimal loss factor for the above definition and $\rho_{s}$ stands for the optimal loss factor with the conventions used in (16) and (17), an application of Stirling’s formula yields $\rho_{s}=\Gamma_{s}^{-1/s}$.
Definition 28 a priori depends on the parameters $(a,b,c,A,s)$. We claim that actually it does not depend on $a,b,c,A$ when $s$ is fixed, in the sense that changing $a,b,c$ and $A$ does not affect the loss factor $\Gamma$.
@@
Fix $s>1$ and assume that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss $\Gamma$. Then, the interpolation problem associated with $(a^{\prime},b^{\prime},c^{\prime},A^{\prime},s)$ is solvable with the same loss factor $\Gamma$.
@@
Assume that the interpolation problem associated with $(a,b,c,A,s)$ is solvable with loss factor $\Gamma$. We show that for every $(a^{\prime},b^{\prime},c^{\prime},A^{\prime},s)$ as above, the associated problem is solvable with loss $\Gamma$. Clearly, for any translation $\tau\in\mathbb{R}$, the problem associated with $(a+\tau,b+\tau,c+\tau,A,s)$ is solvable with the same loss factor. We may therefore assume that $a=a^{\prime}=0$. Let us then use a dilatation: if $\varphi\in C^{\infty}[0,b]$ is such that
$|\varphi^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns},$
@@
$\varphi_{\lambda}\in C^{\infty}[0,b/\lambda],\quad\varphi_{\lambda}^{(n)}(t)=\lambda^{n}\varphi^{(n)}(t),\quad|\varphi_{\lambda}^{(n)}(t)|\leq C^{\prime}(\Gamma A\lambda+\epsilon\lambda)^{n}n^{ns}.$
Thus the problem of parameters $(0,b/\lambda,c/\lambda,A\lambda,s)$ is solved with loss $\Gamma$. We can therefore assume that $b=b^{\prime}=1$.
We then claim that if the problem of parameters $(0,1,c,A,s)$ is solvable with loss $\Gamma$ for $0<c<1$, then for every $c^{\prime}\in(0,1)$ the problem of parameters $(0,1,c^{\prime},A,s)$ is solvable with loss $\Gamma$. For, let $\epsilon>0$ and $(a_{n})$ be any sequence such that
@@
$\varphi^{(n)}(c)=a_{n},\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns}.$
@@
$|\psi^{(n)}(t)|\leq C^{\prime\prime}(\Gamma A+\epsilon)^{n}n^{ns},$
(see [reachable_martin, Lemma 3.7]). Moreover, since $\chi$ is 0 outside of $(c^{\prime}-c,1+c^{\prime}-c)$, $\psi$ is well-defined on $\mathbb{R}$ (so it is also defined on $[0,1]$), and has the same derivatives at $c^{\prime}$ than $\varphi$ has at $c$. Thus the problem of parameters $(0,1,c^{\prime},A,s)$ is solvable, with the same loss $\Gamma$.
Next we show that the cases $c=0$ or $c=1$ are equivalent. Assume for instance that the problem of parameters $(0,1,0,A,s)$ is solvable with loss $\Gamma$. For any sequence $(a_{n})$ such that
@@
$\varphi^{(n)}(c)=(-1)^{n}a_{n},\quad|\varphi^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns}.$
@@
Assume now that the problem of parameters $(0,1,0,A,s)$ is solvable with loss $\Gamma$. We show that, for every $0<c<1$ the problem of parameters $(0,1,c,A,s)$ is solvable with loss $\Gamma$. In fact we will show that (1,1,0,A,s)(-1,1,0,A,s) is solvable with loss $\Gamma$, which is easily seen as a stronger statement, by translation and dilatation arguments. We let $(a_{n})$ be a sequence with
@@
We introduce the “odd” and “even” sequences
@@
$|\varphi^{(n)}(t)|,|\psi^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns},\quad\varphi^{(n)}(0)=a_{n}^{o},\quad\psi^{(n)}(0)=a_{n}^{e}.$
@@
$|\tilde{\psi}^{(n)}(t)|,|\tilde{\varphi}^{(n)}(t)|\leq C^{\prime}(\Gamma A+\epsilon)^{n}n^{ns}.$
@@
We can thus put $c=c^{\prime}=0$, hence we are left to show that if the problem $(0,1,0,A,s)$ is solvable with loss $\Gamma$, then, for every $A^{\prime}>0$, the problem $(0,1,0,A^{\prime},s)$ is solvable with loss $\Gamma$. Assume that the problem associated with $(0,1,0,A,s)$ is solvable with loss $\Gamma$, using the dilatation $\varphi_{\lambda}$ we see that for every $\lambda>0$, the problem associated with $(0,1/\lambda,0,A\lambda,s)$ is solvable with loss $\Gamma$. We distinguish two cases: if $\lambda\leq 1$, then the problem associated with $(0,1,0,\Gamma,A\lambda,s)$ is trivially solvable. If on the contrary $\lambda>1$, using a cutoff that is Gevrey of order $1<r<s$, constant to $1$ near 0 and flat at $1/\lambda$, we see that the problem associated with $(0,1,0,\Gamma,A\lambda,s)$ is also solvable with loss $\Gamma$, as in the case of the equivalence for the parameter $c$. Since $\lambda>0$ is arbitrary, this finishes the proof. ∎
Therefore, the only relevant numbers are $\Gamma$ and $s$.
@@ A.2 A proof of Mitjagin’s Theorem
    We have the estimation of the Taylor coefficients of $\Phi$
@@
    where recall that $r=|s|$. In view of the definition of the $\|\cdot\|_{C^{\infty}_{\beta B}}$-norm, the estimate (49) thus boils down to proving that
@@
    i.e. it is of order $1$ and of type 0 on $\mathbb{C}$. Note however that on the real axis we have a better estimate, of the form
@@
    Let $0<\rho<1$ and $M,K\geq 0$. Assume that $g\in\mathcal{H}(\mathbb{C}_{+})\cap C^{0}(\mathbb{C}_{+}\cup i\mathbb{R})$ is of order 1 with minimal type 0, together with
@@
    Let us show this version of the Phragmén-Lindelöf principle. We will use the following “standard” version of the Phragmén-Lindelöf principle, see [K1, §III.C].
    Let $\alpha\geq 1/2$, $\Omega=\{re^{i\theta}:r>0,\quad|\theta|<\frac{\pi}{2\alpha}\}$, and $g\in\mathcal{H}(\Omega)\cap C^{0}(\overline{\Omega})$ of order $\rho<\alpha$, such that
@@
    $g_{\kappa}(z)=\exp\left(-\frac{K}{\cos(\pi\rho/2)}z^{\rho}-\kappa z\right)g(z)\in\mathcal{H}(\mathbb{C}_{+})\cap C^{0}(\mathbb{C}_{+}\cup i\mathbb{R}).$
    We apply the “standard” Phragmén-Lindelöf’s method quoted above to the function $g_{\kappa}$ on the domain $\Omega:=\{x+iy:x,y>0\}$. Note that the quoted result is invariant by rotation, hence we may see $\Omega$ as a sector with angle $\pi/2$, and the parameter $\alpha$ is set to $2$. On $\Omega$ the function $g_{\kappa}$ is of order $\rho=1$, which is strictly less than $\alpha$. Moreover, on the boundary of $\Omega$ lying on the imaginary axis there holds
@@
    We take $\eta=\kappa$, so that for every $x>0$,
@@
    $|f_{n}|=\left|\frac{\Phi^{(n)}(0)}{n!}\right|\leq\frac{1}{r^{n}}\max_{|s|=r}|\Phi(s)|\leq\frac{1}{r^{n}}\|f\|C_{1}\exp\left[\frac{\beta}{e\cos(\pi/2\beta)}\left(\frac{1+\delta}{B}r\right)^{1/\beta}\right],$
@@
    $|f_{n}|\leq\|f\|C_{1}\left(\frac{e\mu}{n\beta}\right)^{n\beta}=C_{1}\|f\|\left(\frac{1+\delta}{B\cos^{\beta}\frac{\pi}{2\beta}}\right)^{n}n^{-n\beta}.$
@@
where the inclusions are continuous and dense [^6] That the first inclusion is bounded is trivial. That the second inclusion is bounded can be shown using the Sobolev embeding $H^{1}(-1,1)\subset C[-1,1]$. That the inclusions are dense is shown by mollifying with a kernel function lying in $C^{\infty}_{\gamma C}$ for some $1<\gamma<\beta$ and $C>0$.. We will need a similar construction for our spaces of sequences. For a sequence $(w_{n})$ of positive numbers we define $K(w_{n})$ as the set of these sequences $a=(a_{n})$ such that
@@
which is what we aim at showing until the end of the proof. From the density of the set of finitely supported sequences in $Y$, which we denote by $\mathbb{C}^{(\mathbb{N})}$, it is enough to show (56) when $a\in\mathbb{C}^{(\mathbb{N})}$. For such $a$ consider the concentrated functional
$f:=\sum_{n=0}^{\infty}a_{n}\delta^{(n)},\quad\langle\delta^{(n)},\phi\rangle:=\phi^{(n)}(0).$
@@
$\sum_{n=0}^{\infty}|B^{n}n^{n\beta}a_{n}|^{2}\leq c\|f\|_{X^{\prime}}^{2},$
with a constant $c>0$ not depending on $a$. We then invoke Lemma 31 for the concentrated functional $f$. To this end, observe that
@@
so that Lemma 31 applies to $f$. Plain computations show that the function $\Phi$ constructed at Lemma 31 satifies
$\Phi(s)=\sum_{n=0}^{\infty}a_{n}(is)^{n},$
@@
$|a_{n}|\leq C\|f\|_{X^{\prime}}\left(\frac{1+\delta}{\Gamma_{\beta}B+\epsilon/2}\Gamma_{\beta}\right)^{n}n^{-n\beta},$
for a constant $C=C(\delta,\beta,B,\epsilon)$, where $\delta>0$ is to be tuned later on. We deduce
$|B^{n}n^{n\beta}a_{n}|\leq C\|f\|_{X^{\prime}}\left(\frac{1+\delta}{1+\frac{\epsilon}{2\Gamma_{\beta}B}}\right)^{n},$
@@
Douglas’s lemma yields a slightly better conclusion than the inclusion (56): there exists a bounded operator $\mathfrak{R}:\ell^{2}_{\beta B}\rightarrow L^{2}_{\beta,\Gamma_{\beta}B+\epsilon}$ such that $i=\mathcal{B}\mathfrak{R}$. In other words, the interpolation problem in the Gevrey class of order $s$ can be solved continuously and linearly. Compare with [mitjagin_bases, Proposition 21, §7, p. 119].